Your preliminary market analysis of two stocks has yielded the table below. The expected rate of return on the market is equal to 10% and the risk-free rate is 2%. Estimated dividends and earnings are for the next year and are expected to grow at a constant rate. The ratio of dividend payouts relative to earnings is expected to remain constant over time.

aWatch

bWatch

Return on New Investments, ROI

14%

12%

Systematic Risk, b

1.5

1

Estimated earnings per share, E1

$5.00

$4.00

Dividend payout ratio, DPR

60%

60%

Current market price per share, P0

$40.00

$42.00

a. What is the intrinsic value of each of the two stocks using a constant-growth dividend discount model? Should you buy or sell the two stocks?

b. The industry including aWatch and bWatch has an average forward-looking price-earnings ratio (P0/E1) of 9.25. What is the intrinsic value of the two stocks using the multiplier method? Should you buy or sell the two stocks?  

c. How should executives in the two companies adjust the dividend distribution policy to maximize shareholder value? Briefly discuss whether the two companies should increase or lower the dividend payout ratios.
Answer #1

Answers

Wn1.Calculation of cost of equity using CAPM approch

Ke= Rf+Beta8(Rm-Rf)

where,

Ke= Cost of equity, Rf= Risk free rate, Rm= Return on market

ke of aWatch= 2+1.50(10-2)

= 2+12

Ke of aWatch = 14%

ke of bWatch= 2+1.00(10-2)

= 2+8

Ke of bWatch = 10%

Wn2. Calculation of DPS and growth rate

Sr. No. Particulars aWatch bWatch
a Earning per share $5.00 $4.00
b Dividend payout ratio 60% 60%
c dividend per share (a*b) $3.00 $2.40
d Return on new investment 14% 12%
e Retention ratio(1-b) 40% 40%
f Growth rate (d*e) 5.60% 4.80%

a. Calculation of intrinsic value of stock using constant dividend growth model

P= D(ke-g)

where,

P= price of stock, D= dividend per share, Ke= cost of equity, g= Growth rate

i)Stock price of aWatch

P= $3.00/(14%-5.60%)

P= $3.00/8.40%

P= $35.71

ii)Stock price of bWatch

P= $2.40/(10%-4.80%)

P= $2.40/5.20%

P= $46.15

Decision regarding buy or sell of stock

Sr. No Particulars aWatch bWatch
a Current market price $40.00 $42.00
b Intrinsic value $35.71 $46.15
c Remark regarding vulue of stock Overvauled Undervalued
d Decision Sell Buy

b. Calculation of value of stock using multiplier approch

Sr. No. Particulars aWatch bWatch
a Price-earning ratio(times) 9.25 9.25
b Earning per share $5.00 $4.00
c Stock price (a*b) $46.25 $37.00
d Current market price $40.00 $42.00
e Remark regarding value of stock Undervalued Overvalued
f Decision Buy Sell

c. Dividend policy of companies

If Return on equity (ROI) is more than cost of equity (Ke), then such firm are known as growth firm as per Walter's dividend policy in order to maximize value of share dividend payout ratio should have zero